# Heavy baryons in the relativized quark model with chromodynamics

### v1

@@ IV Summary
In this work, we have systematically studied the masses of the heavy baryons in the relativized quark model, namely the (modified) Godfrey-Isgur [(M)GI] model. The key ingredients of the GI model are a universal one-gluon-exchange interaction and a linear confinement potential motivated by QCD. The relativistic effects are included semi-quantitatively. In the MGI model, the linear confinement interaction is replaced by a screened one. The parameters are fitted from the ground state baryons, lowest $1/2^{-}$ and $3/2^{-}$ singly heavy baryons and lowest $\frac{3}{2}^{+}\Lambda_{Q}$ states. We found that all heavy baryons observed can be explained as three-quark states. In particular,
@@
    The $\Omega_{c}(3000)$, $\Omega_{c}(3050)$, $\Omega_{c}(3065)$ and $\Omega_{c}(3090)$ states are $p_{\lambda}$ excitations with quantum numbers $1/2^{-}$, $3/2^{-}$, $3/2^{-}$ and $5/2^{-}$, while the $\Omega_{c}(3120)$ is a $3/2^{-}$ state with the $p_{\rho}$ excitation. The higher state $\Omega_{c}(3188)$ is the $2s_{\lambda}$ excitation with quantum numbers $1/2^{+}$, and $\Omega_{c}(3327)$ is a $d_{\lambda}$ excitation with quantum numbers $3/2^{+}$ or $5/2^{+}$.
@@
    The $\Omega_{b}(6316)$, $\Omega_{b}(6330)$, $\Omega_{b}(6340)$ and $\Omega_{b}(6350)$ states are $p_{\lambda}$ excitations with quantum numbers $1/2^{-}$, $3/2^{-}$, $3/2^{-}$ and $5/2^{-}$, while a narrow $p_{\lambda}$ excited state $\Omega_{b}(6446/6457,3/2^{-})$ is predicted to be about 100 MeV above them.
@@
    The $\Sigma_{c}(2800)$ is most likely a $3/2^{-}$ state with the $p_{\lambda}$ excitation, while the $\Sigma_{b}(6097)$ is a $1/2^{-}$ or $3/2^{-}$ state with the $p_{\lambda}$ excitation.
@@
    The quantum numbers of $\Lambda_{c}(2880)$ and $\Lambda_{c}(2940)$ states are known to be $5/2^{+}$ and $3/2^{-}$ experimentally. Our calculation suggests that they are $d_{\lambda}$ and $p_{\rho}$ excitations. The $\Lambda_{c}(2765)$ might be a $1/2^{+}$ or $1/2^{-}$ state, and the $\Lambda_{c}(2910)$ could be a $1/2^{+}$ or $3/2^{+}$ one. Furthermore, the $\Lambda_{b}(6070)$ is a $1/2^{+}$ state of the $2s_{\lambda}$ excitation, while the $\Lambda_{b}(6146)$ and $\Lambda_{b}(6152)$ states might belong to the $\{3/2^{+},5/2^{+}\}$ or $\{1/2^{-},3/2^{-}\}$ doublet.
@@ Acknowledgments
$9/2^{-}$ MGI 6793 6966 7044 7098
Figure 3: Masses of the $\Sigma_{c}$ baryons (in units of MeV). Here, the black/blue lines stand for the results obtained by the GI/MGI models, while the red lines represent the experimental results taken from PDG [15]. The states used for fittings are indicated by daggers ($\dagger$)
Figure 4: Masses of the $\Sigma_{b}$ baryons (in units of MeV). Here, the black/blue lines stand for the results obtained by the GI/MGI models, while the red lines represent the experimental results taken from PDG [15]. The states used for fittings are indicated by daggers ($\dagger$)
Figure 5: Masses of the $\Lambda_{c}$ baryons (in units of MeV). Here, the black/blue lines stand for the results obtained by the GI/MGI models, while the red lines represent the experimental results taken from PDG [15] and Ref. [5]. The states used for fittings are indicated by daggers ($\dagger$)
Figure 6: Masses of the $\Lambda_{b}$ baryons (in units of MeV). Here, the black/blue lines stand for the results obtained by the GI/MGI models, while the red lines represent the experimental results taken from PDG [15] and Ref. [5]. The states used for fittings are indicated by daggers ($\dagger$)
Figure 7: Masses of the $\Xi_{c}$ baryons (in units of MeV). Here, the black/blue lines stand for the results obtained by the GI/MGI models, while the red lines represent the experimental results taken from PDG [15] and Ref. [5]. The states used for fittings are indicated by daggers ($\dagger$)
Figure 8: Masses of the $\Xi_{b}$ baryons (in units of MeV). Here, the black/blue lines stand for the results obtained by the GI/MGI models, while the red lines represent the experimental results taken from PDG [15] and Ref. [5]. The states used for fittings are indicated by daggers ($\dagger$)
Figure 9: Masses of the $\Omega_{c}$ baryons (in units of MeV). Here, the black/blue lines stand for the results obtained by the GI/MGI models, while the red lines represent the experimental results taken from PDG [15] and Ref. [5]. The states used for fittings are indicated by daggers ($\dagger$)
Figure 10: Masses of the $\Omega_{b}$ baryons (in units of MeV). Here, the black/blue lines stand for the results obtained by the GI/MGI models, while the red lines represent the experimental results taken from PDG [15] and Ref. [5]. The states used for fittings are indicated by daggers ($\dagger$)

### v2

@@ IV Summary
In this work, we have systematically studied the masses of the heavy baryons in the relativized quark model, namely the (modified) Godfrey-Isgur [(M)GI] model. The key ingredients of the GI model are a universal one-gluon-exchange interaction and a linear confinement potential motivated by QCD. The relativistic effects are included semi-quantitatively. In the MGI model, the linear confinement interaction is replaced by a screened one. The parameters are fitted from the masses of the ground state baryons, the lowest $1/2^{-}$ and $3/2^{-}$ singly heavy baryons and the lowest $\frac{3}{2}^{+}\Lambda_{Q}$ states. We found that all heavy baryons observed can be explained as three-quark states. In particular,
@@
    The $\Omega_{c}(3000)$, $\Omega_{c}(3050)$, $\Omega_{c}(3065)$ and $\Omega_{c}(3090)$ states are $p_{\lambda}$-dominated excitations with quantum numbers $1/2^{-}$, $3/2^{-}$, $3/2^{-}$ and $5/2^{-}$, while the $\Omega_{c}(3120)$ is a $3/2^{-}$ state dominated the $p_{\rho}$ excitation. The higher state $\Omega_{c}(3188)$ is the $2s_{\lambda}$ excitation with quantum numbers $1/2^{+}$, and $\Omega_{c}(3327)$ is a $d_{\lambda}$ excitation with quantum numbers $3/2^{+}$ or $5/2^{+}$.
@@
    The $\Omega_{b}(6316)$, $\Omega_{b}(6330)$, $\Omega_{b}(6340)$, and $\Omega_{b}(6350)$ states are $p_{\lambda}$-dominated excitations with quantum numbers $1/2^{-}$, $3/2^{-}$, $3/2^{-}$, and $5/2^{-}$, while a narrow $p_{\rho}$-dominated excited state $\Omega_{b}(6446/6457,3/2^{-})$ is predicted to be about 100 MeV above them.
@@
    The $\Sigma_{c}(2800)$ is most likely a $3/2^{-}$ state dominated by the $p_{\lambda}$ excitation, while the $\Sigma_{b}(6097)$ is a $1/2^{-}$ or $3/2^{-}$ state dominated by the $p_{\lambda}$ excitation.
@@
    Our calculation reproduced the masses of the well-known $\Lambda_{c}(2880)$ and $\Lambda_{c}(2940)$ states, whose quantum numbers are $5/2^{+}$ and $3/2^{-}$ respectively. The $\Lambda_{c}(2765)$ might be a $1/2^{+}$ or $1/2^{-}$ state, and the $\Lambda_{c}(2910)$ could be a $1/2^{+}$ or $3/2^{+}$ one. Furthermore, the $\Lambda_{b}(6070)$ is a $1/2^{+}$ state dominated by the $2s_{\lambda}$ excitation, while the $\Lambda_{b}(6146)$ and $\Lambda_{b}(6152)$ states might belong to the $\{3/2^{+},5/2^{+}\}$ or $\{1/2^{-},3/2^{-}\}$ doublet.
@@ Acknowledgments
MGI 6793 6966 7044 7098
Figure 3: Masses of the $\Sigma_{c}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 4: Masses of the $\Sigma_{b}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 5: Masses of the $\Lambda_{c}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15] and Ref. [5]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 6: Masses of the $\Lambda_{b}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 7: Masses of the $\Xi_{c}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 8: Masses of the $\Xi_{b}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 9: Masses of the $\Omega_{c}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).
Figure 10: Masses of the $\Omega_{b}$ baryons (in units of MeV). Here, the black/blue lines stand for the numerical results obtained by the GI/MGI models. The red lines represent the experimental results taken from PDG [15]. In the first column, we list the experimental observed states whose quantum numbers are not determined yet. The states used for fittings are indicated by daggers ($\dagger$).